OpenStudy (lauluisa):

Write the equation of the line in slope-intercept form and identify the slope -5x+10y=20

Nnesha (nnesha):

slope intercept form equation is y = mx + b you have to solve for y so first thing you have to move -5x on right side

Nnesha (nnesha):

5x is negative on left side to move you have to do opposite so add 5x both side  let me know what you get after this

Nnesha (nnesha):

-5x +5x = ??

Nnesha (nnesha):

you cannot combine 20 and 5x because 5 have variable and 20 doesn't

OpenStudy (lauluisa):

10y=20+5x

OpenStudy (lauluisa):

Then u divide10 by 20 and 5... and its going to be.. y= 1+ 5/10

Nnesha (nnesha):

 20/10= ??? and remember 5 have x so when you divide 10 it should be 5/10x so you can reduce this